A man is twice as old as his son. Ten years ago, he was thrice as old as the son. what is the man's age today?

Answer by LinnW(1048)

Set M = man's age
S = son's age
M = 2S
M - 10 = 3(S - 10)
Substitute 2S for M in M - 10 = 3(S - 10)
2S - 10 = 3(S - 10)
2S - 10 = 3S - 30
add -2S to each side
- 10 = S - 30
add 30 to each side
20 = S
Since M = 2S
M = 2(20) = 40